The human MN blood type is determined by two codominant alleles, LM and LN. The frequency of LM on a small island is 0.80. a. If random mating takes place in this population, what are the expected frequencies of the MM, MN, and NN blood types on the isla

please solved this question
Answer #1

answer: According question expected frequencies of the M, MN, and N blood types on the island

f (LM) = p = 0.80, then f (LN) = q = 1-p = (1 - 0.80) = 0.20

frequency of (LM LM) = p2 + Fpq = (0.80)2 + 0.05 (0.80) (0.20) = 0.64 + 0.008 = 0.648

frequency of (LM LN) = 2pq - 2 Fpq = 2 (0.80) (0.20) - 2 (0.05) (0.80) (0.20) = 0.2 - 0.016 = 0.304

frequency of ( LN LN) = q 2+ Fpq = (0.20)2 + 0.05 (0.80) (0.20) = 0.04 + 0.008 = 0.048


answered by: ANURANJAN SARSAM